### v4

# Zhuoming Chen    Ranajoy Sadhukhan    Zihao Ye    Yang Zhou    Jianyu Zhang    Niklas Nolte    Yuandong Tian    Matthijs Douze    Leon Bottou    Zhihao Jia    Beidi Chen [ [ [ [

Abstract
Large language models (LLMs) with long context windows have gained significant attention. However, the KV cache, stored to avoid re-computation, becomes a bottleneck. Various dynamic sparse or TopK-based attention approximation methods have been proposed to leverage the common insight that attention is sparse. In this paper, we first show that TopK attention itself suffers from quality degradation in certain downstream tasks because attention is not always as sparse as expected. Rather than selecting the keys and values with the highest attention scores, sampling with theoretical guarantees can provide a better estimation for attention output. To make the sampling-based approximation practical in LLM generation, we propose MagicPIG, a heterogeneous system based on Locality Sensitive Hashing (LSH). MagicPIG significantly reduces the workload of attention computation while preserving high accuracy for diverse tasks. MagicPIG stores the LSH hash tables and runs the attention computation on the CPU, which allows it to serve longer contexts and larger batch sizes with high approximation accuracy. MagicPIG can improve decoding throughput by up to 5×5$\times$5 × across various GPU hardware and achieve 54ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct model with a context of 96k tokens.
1 Introduction
Figure 1: While $\mathrm{TopK}$ attention performs well on retrieval tasks (niah) where the useful information reduces to a few words, it degrades severely in aggregated tasks like word extraction (cwe, fwe). x-axis: proportion of attention keys used for $\mathrm{TopK}$ attention.
Large language models (LLMs) with long context windows, such as GPT (Achiam et al., 2023), Llama (Dubey et al., 2024), and Gemini (Team et al., 2023), have gained significant attention for their ability to enhance applications like chatbots (Chiang et al., 2024), search engines (Wang et al., 2024), and video analysis (Cheng et al., 2024). However, serving long-context LLMs is highly challenging due to the unique bottleneck in auto-regressive generation—the key-value (KV) cache, which stores intermediate attention keys and values to avoid re-computation (Pope et al., 2022; Zhang et al., 2023b). Specifically, the KV cache grows linearly with both the batch size and sequence length, occupying substantial GPU memory and increasing decoding time. Moreover, the KV cache makes LLM generation extremely memory-bound, leading to underutilization of GPU computational power. For instance, an NVIDIA A100-40GB GPU can only handle a single request for Llama with a 128k context length, with nearly half of the decoding time spent accessing the KV cache, and poor GPU utilization  (He and Zhai, 2024).
Leveraging the common insight that attention is naturally sparse, dynamic sparse or $\mathrm{TopK}$-based approximation has been extensively studied (Tang et al., 2024; Singhania et al., 2024; Zhang et al., 2024; Wu et al., 2024), but three major challenges prevent a wide adoption in LLM serving systems. (1) Quality Degradation. They usually propose various strategies to approximate a subset of KV cache that yields the highest attention scores. However, $\mathrm{TopK}$ attention itself is a biased attention approximation and lacks theoretical guarantees. Figure 1 shows that even exact $\mathrm{TopK}$ attention results significantly degrade the accuracy of certain downstream tasks. (2) High Overhead. There is a large overhead to identify $\mathrm{TopK}$ attention, which becomes the bottleneck rather than the attention computation. For example, as studied in Liu et al. (2024a), naively applying a search algorithm like IVF (Douze et al., 2024) requires access over $30\%$ key states to obtain the exact $\mathrm{TopK}$, showing an unsatisfying trade-off between search accuracy and cost. (3) No Memory Saving. Although saving KV cache loading time, they cannot reduce the total memory occupied by the KV cache, which limits the maximum context and batch sizes when VRAM is scarce.
An ideal sparse attention approximation approach should (1) preserve full accuracy for a diverse set of downstream tasks with guarantees, (2) involve low-cost overhead for KV cache selection, and (3) save GPU memory. The following observations, together with the performance drop shown in Figure 1 suggest that to achieve such demanding requirements, we need to go beyond $\mathrm{TopK}$ attention:
[…]
  • Seemingly high sparsity is usually a consequence of an attention sink. Most of the attention scores concentrate on initial tokens (attention sink phenomenon) (Xiao et al., 2023), making the distribution look sparser. However, as shown in Figure 2(b), attention scores are distributed more uniformly among tokens except for the sink. According to the geometrical interpretation of sink, keys, and queries shown in  Figure 2(c), the attention sink, which we found surprisingly almost static regardless of the input token, is just for imposing sparsity on the attention distribution.
  • It is hard to find $\mathrm{TopK}$ attention. Figure 2(c) also shows why searching for the Top-K keys is intrinsically costly. The keys and queries usually lie within two narrow cones with nearly opposite orientations, except for the attention sink. This significant mismatch between query and data distributions causes nearest-neighbor search methods to perform poorly.
Figure 2: Left: Examples of long-tailed distribution in LLM. The x-axis is the fraction (or number of tokens) used in the $\mathrm{TopK}$, a.k.a. the sampling budget. Mid: Sink tokens make attention score look sparser. Right: The geometry of attention. The key of attention sink $k_{sink}$ is almost opposite to other tokens, and its orientation is surprisingly invariant with input tokens. Query states lie close to $k_{0}$, thus forming attention sink and Figure 2(b). $k$ usually lies in a narrow cone that is far away from $q$. In certain heads, this geometry will result in a long-tailed distribution of attention score and difficulty searching for the $\mathrm{TopK}$ keys.
These limitations of $\mathrm{TopK}$ attention require rethinking the sparse attention approximation. Rather than only using the keys and values with the highest scores, leveraging information on the distribution can make the estimation more accurate. We approach this as a bias correction problem in sampling. Unbiased and efficient sampling has been long studied in biology (Lukacs, 2009), sociology (Chen et al., 2018) as well as machine learning (Backurs et al., 2019; Chen et al., 2019; Zandieh et al., 2023), with theoretical guarantees.
Figure 3 shows that sampling values according to their corresponding attention score (we call this oracle sampling) achieves a much lower (up to 4×4$\times$4 ×) estimation error than the naive $\mathrm{TopK}$ selection. Deploying sampling estimation in attention is promising, but three challenges remain. First, how a reduction of the attention error can make a difference in downstream performance is unclear (Backurs et al., 2019, 2018). Second, modeling the attention score distribution is necessary for efficient sampling, but inferring the distribution parameters requires expensive computations. Third, fully leveraging the resources of modern hardware, GPU and CPU, with a theoretically efficient algorithm is non-trivial.
Figure 3: $\mathrm{TopK}$ v.s. Sampling, 16k total context
This paper proposes Magic samPlIng for Generation (MagicPIG), which leverages Locality sensitive hashing (LSH) sampling for efficient LLM generation. LSH is employed for sampling to approximate the attention score distribution and estimate attention output. By computing hash functions on GPU and conducting sampling on CPU, MagicPIG can allow massive hash tables and hash functions compared to prior work (Kitaev et al., 2020; Chen et al., 2021), which are of vital importance for accurate estimation (Backurs et al., 2018). Following the practice of Aminabadi et al. (2022); He and Zhai (2024), we offload the KV cache computation, which is memory bound, to CPU to allow a larger batch or longer context. Specifically,
  • In Section 3, we analyze the failures of $\mathrm{TopK}$ attention. Moreover, we study sampling-based attention estimation assuming an oracle for the key distribution (Oracle Sampling Estimation) and empirically demonstrate that it is consistently more effective both for distribution estimation and downstream tasks.
  • In Sections 4.1, 4.2 and 4.3, we present a sampling algorithm to approximate oracle sampling for attention estimation based on locality sensitive hashing and the intuition and motivation from statistic perspectives. To our best knowledge, MagicPIG is the first to leverage LSH sampling in self-attention in decoder-only LLM generation.
  • In Section 4.4, we present our system design to efficiently offload attention computation on the CPU, breaking the memory limit of the GPU for serving larger batches or longer contexts. We also overcome the new challenges of computation and memory size raised by our sampling algorithm to support a larger scale of hashing tables beyond prior work (Chen et al., 2021; Kitaev et al., 2020).
In Section 5, we show the empirical evaluation results of the performance of MagicPIG, demonstrating the accuracy and efficiency. While maintaining high accuracy for diverse tasks, MagicPIG can improve serving throughput by 1.55×1.5$\sim$ 5$\times$1.5 ∼ 5 × (A100, L20, RTX 4090) and can achieve 54ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct  (Dubey et al., 2024) with 96K context. More importantly, we show that MagicPIG already outperforms $\mathrm{TopK}$ attention in the two aggregation tasks in Figure 1, suggesting that sampling indeed goes beyond $\mathrm{TopK}$ attention.
2 Background
In this section, we formulate the targeted attention estimation problem and related works.
2.1 Problem formulation
In LLM decoding phase, self-attention part calculates a weighted average of previous values by
$o=\mathrm{Softmax}(\frac{qK^{T}}{\sqrt{d}})V=wV\quad q\in\mathbb{R}^{1\times d}\quad K,V\in\mathbb{R}^{n\times d}\quad w\in\mathbb{R}^{1\times n}$ (1)
where $d$ is the head dimension and $n$ is the context size. $K=[k_{1},k_{2},...,k_{n}],V=[v_{1},v_{2},...,v_{n}],k_{i},v_{i}\in\mathbb{R}^{1\times d}$ is KV cache. Normalized attention weight $w=\mathrm{Softmax}(\frac{qK^{T}}{\sqrt{d}})\in\mathbb{R}^{1\times n}$ is also called attention (score) distribution. Our target is to find sampling matrix $\Pi\in\mathbb{R}^{n\times m}$ and diagonal matrix $D\in\mathbb{R}^{m\times m}$ which minimize
$\delta=||wV-w\Pi D\Pi^{T}V||$ (2)
where $m\ll n$ is computation budget. For $\mathrm{TopK}$ attention, suppose $w_{r_{1}}>...>w_{r_{m}}>...>w_{r_{n}}$, then
$\Pi_{i,j}=\left\{\begin{aligned} 1&,&\text{if }i=r_{j},\\ 0&,&\text{otherwise}.\end{aligned}\right.\quad D_{ii}=\frac{1}{\sum_{i=1}^{m}w_{r_{i}}}$ (3)
2.2 Related works
Efficient Attention. Attention approximation has been long studied. Reformer (Kitaev et al., 2020), KDEformer (Zandieh et al., 2023) and ScatterBrain (Chen et al., 2021) tackle the problem via locality sensitive hashing. These methods work in training and encoder models like BigGAN (Brock et al., 2019). Theoretically, the error bounds and minimal workload required are continuously improved (Brand et al., 2023; Alman and Song, 2023) but have not proven to be practical for wall-clock acceleration in LLM decoding. Besides, flash-attention (Dao et al., 2022b; Dao, 2023; Dao et al., 2022a), flash-decoding (Ye et al., 2024; Hong et al., 2024) and SlimAttention (He et al., 2024) losslessly accelerate scaled product attention operator by maximizing the utilization of hardware, which is orthogonal to our approach.
Locality sensitive hashing. Locality sensitive hashing (LSH) (Backurs et al., 2019, 2018) is a family of hashing functions which assigns the same hash codes for similar inputs with higher probability than others (Chen et al., 2020b; Jafari et al., 2021). LSH uses two hyper-parameters, $(K,L)$. $L$ hash tables are independently built. Each hash table has its own function $H$ which projects a high-dimension vector to an integer by concatenating $K$ random independent hash functions. In the sampling process, all vectors that share hash codes in at least one hash table with a query will be collected. SimHash (Charikar, 2002) is the LSH family based on cosine similarity. For a vector $x\in\mathbb{R}^{d}$, SimHash generates a random hyperplane $w$ and returns $\mathrm{Sign}(w^{T}x)$. Vectors share the same sign if and only if the random projection is not in between them. For a random projection, all angles are equally likely, thus the probability that two vectors $x$, $y$ share the same sign is $p=1-\frac{\theta}{\pi}$, where $\theta=\arccos{\frac{xy^{T}}{||x||\cdot||y||}}$. If we have $L$ hash tables each with $K$ random hash functions, the probability of $y$ to be retrieved by query $x$ is $1-(1-p^{K})^{L}$.
KV Cache reduction. To get rid of memory bound introduced by KV cache thus enabling a larger batch size or serving a longer prompt, many methods are proposed to reduce the volume of KV cache. For example, H2(Zhang et al., 2023b), SnapKV (Li et al., 2024) and Keyformer (Adnan et al., 2024) calculate heuristics during the prefilling phase to decide which tokens to preserve for decoding phase. Quest (Tang et al., 2024) and Loki (Singhania et al., 2024) do not evict KV cache but apply dynamic sparsity to reduce KV Cache loading at inference time. Besides the reduction along the dimension of sequence length, methods like KIVI (Liu et al., 2024b) and QServe (Lin et al., 2024) reduce the size of KV Cache by quantization.
3 Rethinking attention sparsity
In this section, we examine $\mathrm{TopK}$ attention, which is the theoretical upper bound of prior search-based algorithms, including both static methods (Zhang et al., 2023b; Li et al., 2024) and dynamic methods (Tang et al., 2024; Singhania et al., 2024; Mao et al., 2024). We show that $\mathrm{TopK}$ is sub-optimal and present another attention approximation based on sampling and estimation with an oracle that improves the accuracy and/or the computation cost.
3.1 Achilles’ heel of TopK attention
Figure 4: $\mathrm{TopK}$ estimation error for a KV-cache of 16k tokens.
As it is defined, $\mathrm{TopK}$ attention only computes the weighted average on elements with the highest attention scores. To quantify its performance, the computation budget of $\mathrm{TopK}$ attention is defined as the number of selected tokens, i.e., the $\mathrm{K}$ of $\mathrm{TopK}$. Searching-based sparse attention algorithms, like (Tang et al., 2024; Singhania et al., 2024; Wu et al., 2024), are approximations for $\mathrm{TopK}$ attention by replacing the true $\mathrm{TopK}$ keys with the ones found by approximate searching algorithms.
However, we find significant performance degradation in downstream tasks caused by $\mathrm{TopK}$ attention as shown in Figure 1. Although $\mathrm{TopK}$ attention preserves accuracy for retrieval tasks that only require a minimal subset of the context (needle-in-a-haystack single/multikey (Hsieh et al., 2024)), it severely degrades for aggregation tasks that leverage the full context (common word extraction and frequent word extraction (Hsieh et al., 2024)). Intuitively, the information is distributed more broadly for aggregation tasks, which results in less peak attention score distribution.
[…]
To better understand the attention distribution, we study the geometry of $q,k$ and make the following three observations. (1) Key states of the initial token (also known as attention sink, denoted by $k_{sink}$) remain almost the same for arbitrary input. In Figure 5(a), we randomly draw $32$ samples from the vocabulary and measure the mutual cosine similarity of key states. Surprisingly, we find that the orientations of the key states of different input tokens are almost identical with a similarity $>0.99$. (2) The orientation of the center of key states (i.e. $k_{avg}=\frac{1}{n}\sum_{i=1}^{n}k_{i}$) remains stable for different input sentences. In Figure 5(b), we measure the mutual cosine similarity of $k_{avg}$ of $50$ different input sentences. Although variance exists, the similarity of $k_{avg}$ is over $0.9$. (3) The orientations of $k_{avg}$ and $k_{sink}$ are almost opposite. In Figure 5(c), we find that for each head, $k_{sink}$ and $k_{avg}$ has a cosine similarity between $-0.9\sim-0.8$.
These observations shape the geometry as shown in Figure 2(c). The attention sink, which is static regardless of input, produces high sparsity in the attention distribution, whereas other parts are more uniformly distributed. Simply applying $\mathrm{TopK}$ will place even more weight on the sink token, thus losing contextual information. In addition, misaligning $q$ and $k$ also causes difficulty in search (Liu et al., 2024a).
Figure 5: Geometric information of attention. Left: With arbitrary input, the orientation of $k_{sink}$ almost remains the same, with a minimum similarity $>0.99$ across sampled inputs. Mid: The orientation of $k_{avg}$ is stable across various input sentences with a similarity $>0.9$ observed. Right: $k_{sink}$ and $k_{avg}$ are almost opposite with similarity between $-0.9\sim-0.8$.
3.2 Estimate attention with sampling
Figure 6: Left and Middle: Oracle sampling estimation can significantly reduce numerical error compared to $\mathrm{TopK}$ attention. The evaluated context size is 16k. The $x$-axis is sampling budget for oracle sampling and computation budget for $\mathrm{TopK}$ attention. Notice that the estimation error of $\mathrm{TopK}$ attention will cross oracle sampling after a certain large budget (12k in figures). This is because oracle sampling will repetitively sample the same subset of tokens with a high probability while $\mathrm{TopK}$ will not. Theorem 3.3 further explains this. Right: Downstream comparison for oracle sampling estimation and $\mathrm{TopK}$ attention. The $x$-axis for both methods is computation budget ratio, i.e. the fraction of selected/sampled tokens.
Existing TopK attention mechanisms ignore tokens in the KV cache with low attention scores, which introduces a bias since the ignored tokens comprise a large proportion of attention scores (Figure 2(a)). As a result, TopK attention achieves suboptimal performance for long-context tasks, such as information aggregation (Figure 1). Increasing the computation budget for TopK attention does help reduce the estimation error (Figure 4) since it will involve more elements in computing. However, the following question is posed:
Can we improve the estimation quality with low computational budgets?
Inspired by mark and recapture (Lukacs, 2009; Owen, 2013; Lohr, 2021; Chen et al., 2018), we show in the following that attention output can be estimated with sampling. Using notations from Section 2.1 we can re-write attention output $o$ as the expectation of $v_{i},1\leq i\leq n$ from distribution $w$, i.e. $o=\mathbb{E}_{i\sim w}(v_{i})$, which can be estimated by the following method.
Definition 3.1 (Oracle Sampling Estimation).
Given a sampling budget $\mathcal{B}$ and normalized attention score $w$, $\mathcal{B}$ elements are sampled independently from $w$ (i.e. $i_{1},i_{2},...,i_{\mathcal{B}}\overset{\mathrm{iid}}{\sim}w$). Then the attention output is estimated as
$\bar{o}=\frac{1}{\mathcal{B}}\sum_{j=1}^{\mathcal{B}}v_{i_{j}}$ (4)
This is not the lowest variance estimator but has better downstream performance (see Section 8). We call it “oracle” because it assumes that the exact attention vector $w$ is known, which is not true for sparse attention approximations.
Theorem 3.2.
Oracle sampling estimation is unbiased, and the trace of covariance monotonically decreases with $\mathcal{B}$.
This theorem (proved in Section 7) theoretically guarantees a low estimation error of oracle sampling. We also present an empirical comparison between oracle sampling estimation and $\mathrm{TopK}$ attention in Figures 6(a) and 6(b). In summary, oracle sampling estimation can reduce relative error by up to 4×4$\times$4 ×.
Note that the sampling budget $\mathcal{B}$ is not the actual computation cost for oracle sampling estimation: duplicate $X_{i}$ need to be computed/loaded only once, so $\bar{o}$ can be computed by
$\bar{o}=\sum_{i\in\mathcal{S}}\frac{f_{i}}{\mathcal{B}}v_{i}\quad{S}=\mathrm{Unique}(\{i_{1\leq i\leq\mathcal{B}}\})$ (5)
where $f_{i}$ is the number of duplicates of $X_{i}$. Intuitively, if $w$ has an peaked distribution (e.g. $w_{i}>99\%$), then almost all samples in $\{i_{1},...,i_{\mathcal{B}}\}$ are identical to $i$. The actual computation cost of oracle sampling estimation is $|{S}|$, the number of unique samples, which we bound in the following:
Theorem 3.3.
The expected computation budget ($\mathbb{E}({|{S}|}))$ has an upper bound of $1+\mathcal{B}\epsilon$, where $\epsilon=1-\max_{i}w_{i}$.
This theorem (proved in Section 7) shows that the computation cost of oracle sampling is usually far less than the sampling budget. In Figure 6(c), we present the downstream accuracy comparison between oracle sampling estimation and $\mathrm{TopK}$ attention. The former preserves high accuracy for both tasks, even with a very small computation cost ($0.002\%$ out of 16k context, which is approximately $32$). In Section 12, we provide an intuitive example to explain why sampling outperforms $\mathrm{TopK}$ in estimation.
4 MagicPIG
Section 3.2 demonstrates the potential of sampling-based estimation. In Sections 4.1 and 4.2, we present how we arrive at Locality sensitive hashing to unleash this potential from a statistical perspective. In Section 4.3, we show the practical algorithm. Finally, in Section 4.4, we demonstrate our system co-design for accurate and efficient LLM decoding through GPU-CPU collaboration.
Note that most of the derivations in this section might be classical and can even be found in textbooks, but our goal is to leverage them to motivate MagicPIG design and precisely demonstrate the power of a rigorously sound algorithm with system co-design in deep generative models.
4.1 Self-normalized importance sampling for attention estimation
Oracle sampling estimation cannot go beyond 2×2$\times$2 × wall clock speed up because obtaining distribution $w$ requires full computation of all $qk_{i}^{T}$, thereby only saving the $wV$ computation.
Fortunately, importance sampling (Kloek and Van Dijk, 1978; Owen, 2013; Lohr, 2021) allows us to estimate unknown distribution $w$ by sampling from a proposed distribution $u$. In our problem setting, the normalization factor of $w$, i.e. $Z=\sum_{i=1}^{n}\exp{\frac{qk_{i}^{T}}{\sqrt{d}}}$ is also unknown because computing it requires evaluating all $qk_{i}^{T}$. However, we do have access to unnormalized weights $\widetilde{w_{i}}=e^{\frac{qk_{i}^{T}}{\sqrt{d}}}$ for sampled indices $i$. Hence, by employing a variant of importance sampling, self-normalized importance sampling (Owen, 2013), we sample indices $i_{1},i_{2},...,i_{\mathcal{B}}$ from a proposed distribution $u$ and the resulting estimator is
$X^{\mathrm{IS}}=\frac{1}{\widetilde{Z}}\sum_{j=1}^{\mathcal{B}}\frac{\widetilde{w_{i_{j}}}}{u_{i_{j}}}v_{i_{j}}\quad\text{where}\quad\widetilde{Z}=\sum_{j=1}^{\mathcal{B}}\frac{\widetilde{w_{i_{j}}}}{u_{i_{j}}}$ (6)
which has a very nice property for accurately estimating attention output that $\mathbb{P}[\lim_{\mathcal{B}\rightarrow\infty}X^{\mathrm{IS}}=o]=1$.
Its variance [^1] We assume head dimension $d=1$ here for simplicity. Higher dimensions have similar formulations and analyses by replacing variance with the trace of covariance. is related to the distribution $u$, and can be approximated by
$\widetilde{\mathrm{Var}}(X^{\mathrm{IS}})=\frac{1}{\mathcal{B}}\mathbb{E}_{i\sim u}[\frac{w_{i}^{2}}{u_{i}^{2}}(v_{i}-o)^{2}]=\frac{1}{\mathcal{B}Z^{2}}\mathbb{E}_{i\sim u}[\frac{\widetilde{w_{i}}^{2}}{u_{i}^{2}}(v_{i}-o)^{2}]$ (7)
To minimize the variance, $u$ should satisfy $u\propto\widetilde{w_{i}}|v_{i}-o|$ (Hesterberg, 2003). The variance will be high if $u_{i}$ and $\widetilde{w_{i}}|v_{i}-o|$ assign a high probability mass to different regions of the sample space or have different modes. Therefore, the challenge is computing a distribution $u$ aligned with $\widetilde{w_{i}}|v_{i}-o|$ without accessing too many $\widetilde{w_{i}}$. Besides, Equation 6 requires that sampling probability $u$ can be computed and $u_{i}>0$, which is not satisfied by many deterministic approximations like $\mathrm{TopK}$.
4.2 Variance reduction with LSH
We decompose $\widetilde{w_{i}}|v_{i}-o|=\exp({\frac{qk_{i}^{T}}{\sqrt{d}}+\log|v_{i}-o|}$). We observe empirically (Figure 10 in the appendix) that $\log|v_{i}-o|$ does not fluctuate significantly compared to $\frac{qk_{i}^{T}}{\sqrt{d}}$. Hence, we simplify the requirement of $u$ to share the same peaks with $qk_{i}^{T}$. By the following transformation,
$r=\max_{1\leq i\leq n}|k_{i}|\quad\bar{q}=[q,0]\quad\bar{k_{i}}=[k_{i},\sqrt{r^{2}-|k_{i}|^{2}}]$ (8)
we further transfer the inner product $qk_{i}^{T}$ to cosine similarity between $\bar{q}$ and $\bar{k_{i}}$ (which is a common practice in Maximum Inner Product Search (Shrivastava and Li, 2014)).
Inspired by prior work (Spring and Shrivastava, 2017; Chen et al., 2020a), we leverage Locality sensitive hashing-based sampling for this estimation problem. Specifically, leveraging a hash function $h$ in the LSH family that preserves cosine similarity such as SimHash (Sadowski, 2007), we can sample from probability distribution $u_{i}=\mathbb{P}[{h}(q)=h(k_{i})]$ which is monotonic to $\cos{\frac{qk_{i}^{T}}{|q|\cdot|k_{i}|}}$.
4.3 Algorithm Design
To make this estimation practical, MagicPIG is implemented by the following specific design.
Estimator approximation. Self-normalized important sampling Equation 6 requires $i_{1},i_{2},...,i_{k}$ $\mathrm{iid}$ sampled, but the probabilities provided by hashing are not normalized. Hence we adapt our estimator: After obtaining $S$ with probability $u$, MagicPIG computes
$X=\frac{\sum_{i=1}^{n}\frac{\widetilde{w_{i}}}{u_{i}}v_{i}\mathbf{1}_{i\in S}}{\sum_{i=1}^{n}\frac{\widetilde{w_{i}}}{u_{i}}\mathbf{1}_{i\in S}}=\frac{\sum_{i\in S}\frac{\widetilde{w_{i}}}{u_{i}}v_{i}}{\sum_{i\in S}\frac{\widetilde{w_{i}}}{u_{i}}}$ (9)
Hash function selection. MagicPIG leverages SimHash (Sadowski, 2007), that draws with $K\times L$ random vectors. For each of the $L$ hash tables, the ${q}$ and ${k_{i}}$s vectors are projected on $K$ directions, and only the sign of the projection is kept, which yields a $K$-bit hash value. Key ${k_{i}}$ is sampled only if there exist at least two [^2] Empirical results show that requiring hits in two hash tables greatly improves accuracy over standard SimHash. hash tables where ${k_{i}}$ shares the hash value with ${q}$. The corresponding probability is
$u_{i}=\mathbb{P}[\text{${k_{i}}$ is sampled}]=1-(1-p^{K})^{L}-Lp^{K}(1-p^{K})^{L-1}\quad\text{where}\quad p=1-\frac{1}{\pi}\arccos{\frac{{q}{k_{i}}^{T}}{|{q}|\cdot|{k_{i}}|}}$ (10)
Input: $\boldsymbol{K},\boldsymbol{V}\in{R}^{n\times d}$, $\boldsymbol{q}\in{R}^{1\times d}$, random projectors $\boldsymbol{W}\in R^{d\times(K\times L)}$, hash tables $\boldsymbol{HT}$, static KV cache $\boldsymbol{K_{T}},\boldsymbol{V_{T}}\in{R}^{t\times d}$.
Compute hash code for new query
[…]
Query hash tables to sample $\boldsymbol{S}$ in Equation 9
$\boldsymbol{S}$ = $\textbf{Query}(\boldsymbol{HT},\boldsymbol{q}_{\text{code}}),\boldsymbol{K_{S}}=\boldsymbol{K[S]},\boldsymbol{V_{S}}=\boldsymbol{V[S]}$
Compute inner product for $\boldsymbol{q}$ and sampled $\boldsymbol{K}$
$\boldsymbol{w_{S}}=\boldsymbol{q}\boldsymbol{K^{T}_{S}}$, $\boldsymbol{w_{T}}=\boldsymbol{q}\boldsymbol{K^{T}_{T}}$
Compute collision probability for each hash function
$\boldsymbol{p}=1-\frac{1}{\pi}\arccos({{\boldsymbol{w}}/{(||\boldsymbol{q}||\cdot||\boldsymbol{K_{S}}||})})$
Compute sampling probability
$\boldsymbol{u}=1-(1-\boldsymbol{p}^{K})^{L}-L\boldsymbol{p}^{K}(1-\boldsymbol{p}^{K})^{L-1}$
Compute attention output estimation
$\boldsymbol{\bar{o}}=\textbf{Softmax}(\frac{[\boldsymbol{w_{S}},\boldsymbol{w_{T}}]}{\sqrt{d}}-\log([\boldsymbol{u},\boldsymbol{1_{t}}]))[\boldsymbol{V_{S}},\boldsymbol{V_{T}}]$
Return $\boldsymbol{\bar{o}}$
Algorithm 1 MagicPIG Decoding
Data pre-processing. Before building hash tables, MagicPIG centers the $k_{i}$ vectors. As shown in Figure 2(c), keys are almost always concentrated on one side of the queries, except the initial token. In this case, random projections cannot effectively distinguish keys, resulting in uniform sampled probabilities. $\mathrm{Softmax}$ is translation invariant. Centering ($\bar{k_{i}}=k_{i}-\frac{1}{n}\sum_{i=1}^{n}k_{i}$) distributed the keys better and remains computationally equivalent.
Combining Equations 9 and 10 gives a closed form of the MagicPIG attention estimation. Assuming sample set $S$ is obtained with LSH,
$\bar{o}$ $=\sum_{i\in S}\frac{\exp{(\frac{qk_{i}^{T}}{\sqrt{d}}-\log u_{i})}}{\sum_{i\in S}\exp{(\frac{qk_{i}^{T}}{\sqrt{d}}-\log u_{i})}}v_{i}$
$u_{i}$ $=1-(1-p_{i}^{K})^{L}-Lp_{i}^{K}(1-p_{i}^{K})^{L-1}$
$p_{i}$ $=1-\frac{1}{\pi}\arccos{\frac{{q}{k_{i}}^{T}}{|{q}|\cdot|{k_{i}}|}}$ (11)
4.4 System co-design
Figure 7: Left: Memory hierarchy of hardware. GPU VRAM has high bandwidth but is limited. CPU DRAM is sufficient but is relatively slow. The limited bandwidth of PCIE forbids large-scale data transfer. Right: Workload partition of MagicPIG. Linear projections and hash function computation (by random projection) are done on GPU, while sampling with hash tables and attention are done on CPU. The execution order is ①③④②circled-1circled-3circled-4circled-2①③④②①③④② at decoding time.
The memory size of KV cache remains a bottleneck for long-context LLM decoding, especially when GPU VRAM is limited. DRAM on the CPU side offers sufficient memory capacity with $100-200$GB/s bandwidth, which is usually $10-20\%$ of GPU VRAM bandwidth (see Figure 7(a)). Ideally, this gap can be mitigated by 510×5-10$\times$5 - 10 × sparsity. To make CPU DRAM an aggregated memory for GPU, the workload must be partitioned. In our experiments, $K=9$ or 10, and $L$ is a few hundred.
Our system design extends prior work (He and Zhai, 2024; Aminabadi et al., 2022) by splitting LLM decoding into three parts. (1) Parameter computations, i.e., all linear projectors including MLP and $W_{Q},W_{K},W_{V},andW_{O}$ in the self-attention module run on GPU. (2) Attention computation, which involves $o=\mathrm{Softmax}(\frac{qK^{T}}{\sqrt{d}})V$, runs on CPU. (3) Random projections. At generation time, for each $q$, $K\times L$ random projections are conducted to obtain the hash codes. Since all heads can share the same random projectors, the memory overhead is limited (400 KB in our implementation), so this step is compute-bound. Therefore, the projection is placed on GPU. (4) Retrieval. The hash codes of $q$, need to be looked up in $L$ hash tables, which is negligible computationally. However, the pre-built hash tables for $k_{i}$s can occupy considerable memory, making it a better fit for the CPU. With the above partition, we are able to support hash tables with $K$ and $L$ beyond the scale of prior work (Kitaev et al., 2020; Chen et al., 2021; Zandieh et al., 2023) without worrying about computation for hash codes as well as the storage of hash tables.
On-device cache. Sink tokens (the first several tokens) and local tokens are more likely to be sampled according to their high similarity to the query. To further reduce CPU workload, MagicPIG stores these tokens on GPU and does not apply LSH sampling to them. We leverage the recursive attention technique (Ye et al., 2024) to merge the attention output from CPU and GPU.
Our algorithm applies to a single attention head, see Algorithm 1. The details of Encode, Query, as well as the hash table construction, are described in prior work (Sadowski, 2007; Chen et al., 2020b). In Section 11, we discuss the selection of LSH hyper-parameter (K, L).
5 Evaluation
In this section, we aim to demonstrate that MagicPIG can speed up LLM decoding while preserving high accuracy. We first present MagicPIG’s accuracy in downstream tasks, followed by our end-to-end system results showing wall-clock performance.
  • In Section 5.1, we demonstrate MagicPIG preserves high accuracy (less than $2\%$ degradation) across moderate to long context tasks with computation cost $2\%\sim 5\%$ of full attention.
  • In Figure 8, we demonstrate the system performance of MagicPIG, which achieves up to 5×5$\times$5 × throughput improvement and 54ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct with 96K context.
  • In Section 5.3, we verify the effectiveness of centering, which is of vital importance for the success of sampling. Also, we demonstrate that MagicPIG already outperforms $\mathrm{TopK}$ attention in the two aggregation tasks in Figure 1, indicating that sampling indeed goes beyond $\mathrm{TopK}$ attention.
5.1 MagicPIG Preserves Accuracy
We demonstrate that MagicPIG can preserve accuracy in diverse tasks with less than $5\%$ computation.
Setup. Our experiments are based on Llama (AI@Meta, 2024; Dubey et al., 2024; Touvron et al., 2023) models. Three types of tasks are included, which are 3 mid-context comprehensive tasks from lm-eval-harness (Gao et al., 2021) (GSM8K-CoT (Cobbe et al., 2021), MMLU-Flan-Cot-Fewshot (Hendrycks et al., 2020) and COQA (Reddy et al., 2019)), and 6 long context tasks from (Bai et al., 2023) (QASPER (Dasigi et al., 2021), LCC, Repobench-P (Liu et al., 2023), TriviaQA (Joshi et al., 2017), PRE and TREC (Li and Roth, 2002; Hovy et al., 2001)) and 13 synthetic tasks from RULER (Hsieh et al., 2024) (with 50 examples per task).
Baselines. Besides full attention, Quest (Tang et al., 2024) and its variants are used as baselines. In its default setting, Quest uses a “page size” of 16, i.e. 1/16 of the full attention cost. To compare the methods fairly in the low computation budget regime, we also evaluate Quest with page size 32 and 64 and make sure at least one page is selected in every test example. The initial 4 tokens and local 64 (for LongBench (Bai et al., 2023) and RULER (Hsieh et al., 2024)) or 24 (for lm-eval-harness (Gao et al., 2021)) tokens as well as layer-$\{0,16\}$ are statically preserved. We do not use the theoretical transformations in Equation 8 in our implementations, as we do not find them to contribute to accuracy improvements.
Cost. The cost for the attention approximation consists of two parts: $\text{Cost}_{1}$ is the sampling/search cost to obtain $S$ in Equation 11, $\text{Cost}_{2}$ is the attention computation cost, see Equation 11. We report the ratio of the number of FLOPs compared to the full attention computation. For MagicPIG, $\text{Cost}_{1}\simeq 0$ and $\text{Cost}_{2}$ is empirically measured for different LSH hyper-parameters. For Quest with page size $K$, $\text{Cost}_{1}=\frac{1}{K}$ and $\text{Cost}_{2}$ is controlled manually.
Analysis. From Tables 1, 2 and 3, (1) MagicPIG preserves high accuracy (degradation less than $2\%$) for all kinds of tasks, with a computation cost of $2\%\sim 5\%$. (2) Compared with Quest, which also shows reasonable performance on long context tasks, MagicPIG also demonstrates good performance on tasks with moderate context sizes in lm-eval-harness (Gao et al., 2021), indicating a more robust performance in general serving. (3) With LSH sampling, which introduces an order of magnitude lower sampling/searching cost ($\text{Cost}_{1}$), MagicPIG can achieve equivalent or better accuracy with only half of the computation cost.
Table 1: Comprehensive tasks on lm-eval-harness (Gao et al., 2021). MagicPIG significantly outperforms other methods with lower computation. The config (K, L) is a hyper-parameter of LSH for MagicPIG or page size and ratio of selected pages for Quest (Tang et al., 2024). $\text{Cost}_{1}$, $\text{Cost}_{2}$ represents the cost for searching/sampling and sparse attention computation.
Methods Config GSM COQA MMLU Avg. $\text{Cost}_{1}$ $\text{Cost}_{2}$ $\text{Cost}_{total}$.
Llama-2-7b-chat Full 22.4 75.8 49.2 49.1 0.00 1.00 1.00
MagicPIG (10,220) 17.3 76.4 48.6 47.4 0.00 0.04 0.04
MagicPIG (8,90) 18.7 75.0 47.9 47.2 0.00 0.08 0.08
Quest (16,0.05) 13.0 69.4 41.4 41.3 0.06 0.05 0.11
Quest (32,0.1) 15.7 70.2 44.0 43.3 0.03 0.10 0.13
Llama-3.1-8B-Instruct Full 77.6 78.5 65.2 73.7 0.00 1.00 1.00
MagicPIG (10,220) 72.7 78.1 62.7 71.2 0.00 0.03 0.03
MagicPIG (8,90) 71.0 78.0 61.3 70.1 0.00 0.07 0.07
Quest (16,0.05) 57.9 64.6 42.5 55.0 0.06 0.05 0.11
Quest (32,0.1) 64.5 65.0 48.0 59.2 0.03 0.10 0.13
Table 2: Long context tasks on LongBench (Bai et al., 2023). MagicPIG preserves high accuracy with low computation. Config and cost are defined as in Table 1. Code models are only evaluated by Repobench-P and LCC.
Methods Config QaS RbP LCC PrE TrC TrQ Avg. $\text{Cost}_{1}$ $\text{Cost}_{2}$ $\text{Cost}_{total}$.
Llama-3.1-8B-Instruct Full 44.9 52.1 66.8 100.0 71.3 91.8 71.2 0.00 1.00 1.00
MagicPIG (10,150) 43.2 50.2 64.4 100.0 71.3 92.2 70.3 0.00 0.02 0.02
MagicPIG (8,75) 43.5 50.4 67.0 100.0 71.7 91.7 70.7 0.00 0.05 0.05
Quest (16,0.05) 45.7 49.7 64.9 100.0 71.7 91.5 70.6 0.06 0.05 0.11
Quest (32,0.1) 44.4 50.5 65.1 100.0 71.3 91.6 70.5 0.03 0.10 0.13
Code-Llama-13b-16K Full 58.5 74.7 66.6 0.00 1.00 1.00
MagicPIG (10,150) 56.9 74.0 65.5 0.00 0.03 0.03
Quest (16,0.05) 56.4 74.4 65.4 0.06 0.05 0.11
Table 3: Synthesized tasks on RULER (Hsieh et al., 2024). MagicPIG preserves high accuracy with low computation. Config and cost are defined as in Table 1.
Methods Config 16K 32K 64K 96K Avg. $\text{Cost}_{1}$ $\text{Cost}_{2}$ $\text{Cost}_{total}$.
Llama-3.1-8B-Instruct Full 94.2 91.5 86.1 83.0 88.7 0.00 1.00 1.00
MagicPIG (10,150) 91.8 88.9 84.8 80.0 86.4 0.00 0.02 0.02
MagicPIG (9,120) 93.4 90.6 84.7 81.5 87.6 0.00 0.04 0.04
MagicPIG (8,75) 92.9 90.2 84.9 81.7 87.4 0.00 0.05 0.05
Quest (16,0.04) 86.3 85.4 81.9 74.9 82.1 0.06 0.04 0.10
Quest (32,0.06) 84.3 84.0 80.1 74.4 80.7 0.03 0.06 0.09
Quest (64,0.08) 85.2 84.3 77.0 74.2 80.2 0.02 0.08 0.10
MegaBeam-Mistral-7B-512K Full 91.7 88.1 83.5 83.7 86.8 0.00 1.00 1.00
MagicPIG (10,150) 89.8 86.5 81.7 80.7 84.7 0.00 0.02 0.02
MagicPIG (9,120) 90.7 88.5 82.9 82.4 86.1 0.00 0.04 0.04
MagicPIG (8,75) 90.6 86.4 82.8 81.6 85.4 0.00 0.05 0.05
Quest (16,0.04) 83.3 83.2 79.3 78.6 81.1 0.06 0.04 0.10
Quest (32,0.06) 81.5 80.8 76.7 74.4 78.4 0.03 0.06 0.09
Quest (64,0.08) 79.6 77.5 73.8 73.7 76.1 0.02 0.08 0.10
Llama3-8B-Prolong-512K Full 93.5 90.8 85.1 83.5 88.2 0.00 1.00 1.00
MagicPIG (10,150) 88.0 86.4 81.3 78.8 83.6 0.00 0.02 0.02
MagicPIG (10,170) 89.0 88.7 82.8 80.0 85.1 0.00 0.025 0.025
MagicPIG (9,120) 91.4 88.2 82.4 80.4 85.6 0.00 0.04 0.04
MagicPIG (8,75) 91.4 88.6 83.1 80.5 85.9 0.00 0.05 0.05
Quest (16,0.04) 84.9 83.7 78.7 78.6 81.5 0.06 0.04 0.10
5.2 MagicPIG Shows Impressive Efficiency across Various Hardware Settings
We show MagicPIG can bring up to 5×5$\times$5 × wall clock speed up and reduce GPU memory consumption on different models and hardware settings (A100, L20, RTX4090).
Figure 8: We evaluate MagicPIG on three serving scenarios. Left: A100 serves 34B model with 16K context. MagicPIG achieves 1.5×1.5$\times$1.5 × throughput improvement. Mid: L20 serves 13B model with 16K context. MagicPIG achieves 5.0×5.0$\times$5.0 × throughput improvement. Right: Simulated RTX 4090 serves 8B model with 96K context. MagicPIG achieves a latency of 54ms in a single request serving and can improve the throughput of baseline by up to $\times3$. The dashed lines denote the highest throughput of baselines. With KV cache offloading, MagicPIG can fit a much larger batch size compared with full attention on GPU, which contributes to the throughput improvement.
Setup. We evaluate our system performance on 3 serving settings. (1) 80GB GPU (A100) and 34B model (CodeLlama-34B) (Rozière et al., 2024) with 16K contexts; (2) 48GB GPU (L20) and 13B model (CodeLlama-13B) (Rozière et al., 2024) with 16K contexts; (3) 24GB GPU [^3] We simulate 24GB GPU by setting memory limit with L20. As the bandwidth of L20 (864GB/s) is less than RTX 4090 (1TB/s), the real speed of our system should be slightly faster than the simulation. (e.g. RTX 4090) and 8B model (Llama-3.1-8B) (Dubey et al., 2024) with 96K contexts.
Baselines. Our baselines for (1) and (2) are full attention on GPU, and for (3) is full attention on CPU with theoretical estimated bandwidth. Our system’s GPU part is implemented in native Pytorch (Paszke et al., 2019) and the CPU part in FBGEMM (Khudia et al., 2021) in bfloat16 precision. Our CPU is Intel Platinum 8480+ for A100 and Intel 8563C for L20. In the last setting, the CPU bandwidth is estimated at 150GB/s, above the empirical bandwidth we measure when running a group query attention of size 4.
Analysis. In Figures 8(a), 8(c) and 8(b), we demonstrate (1) MagicPIG significantly improves decoding throughput for all three scenarios (A100: 1.5×1.5$\times$1.5 ×, L20: 5.0×5.0$\times$5.0 ×, RTX 4090: $\times3$) and can achieve a latency of 54ms for single request generation with 96K context for RTX 4090. (2) With KV cache offloading, MagicPIG can fit much larger batches than GPU full attention baselines (over 12×12$\times$12 ×). The ablation study of decoding throughput with different LSH hyper-parameters is presented in Table 7.
5.3 Ablation Study
Figure 9: Left: Accuracy comparison for with and without centering. Here we fix $K$ and vary $L$ for the two settings. Mid and Right: Comparison between $\mathrm{TopK}$ attention and MagicPIG. In the two aggregated tasks, sampling-based MagicPIG can even beat the exact $\mathrm{TopK}$ attention. The experiments are done on RULER (Hsieh et al., 2024) with a 16K context size.
In this section, we empirically validate our two previous observations.
Centering is important for good performance. In Section 4.3, we use a translation to center the keys before applying LSH sampling. Empirical results show this to be important for downstream tasks as shown in Figure 9(a). Without centering, the accuracy drops to almost zero in retrieval (NIAH) and degrades to $65\%$ in FWE. We find almost no keys (less than $0.1\%$) can be sampled by the query without centering, as their orientation is almost opposite, as shown in Figure 2(c).
Sampling goes beyond $\mathrm{TopK}$. In Figures 9(b) and 9(c), We compare the performance of MagicPIG and $\mathrm{TopK}$ attention in two aggregated tasks (CWE, FWE) where $\mathrm{TopK}$ attention experiences significant performance degradation (Figure 1). MagicPIG can even beat exact $\mathrm{TopK}$ attention in these two tasks by a margin up to $3\%$ and $8\%$ respectively, demonstrating that sampling improves the ceiling of $\mathrm{TopK}$, which is impossible for a search-only algorithm.
6 Conclusion
In this work, we first present the limitation of $\mathrm{TopK}$ attention approximation for addressing the computational and memory challenges of long-context LLM generation. Then we show oracle sampling can go beyond $\mathrm{TopK}$ and introduce MagicPIG, a novel approach that leverages LSH sampling to approximate the oracle sampling. MagicPIG significantly reduces the workload of attention computation while preserving high accuracy across diverse tasks. MagicPIG relies on LSH sampling and a system co-design that offloads hash tables and reduced attention computation to the CPU. Our experimental results demonstrate that MagicPIG substantially improves throughput and latency across multiple hardware configurations, outperforming traditional $\mathrm{TopK}$ attention mechanisms. The theoretical soundness, robustness, and scalability of MagicPIG open up new opportunities in both attention approximation methods and algorithm-hardware co-design.